OpenStudy (b77w):

How do I solve for x in this problem?: sqrt(2x+3)=x

OpenStudy (souvik):

square both sides first

OpenStudy (b77w):

I understand now! Once I square them, I can just factor them.

OpenStudy (souvik):

yes...!

OpenStudy (anonymous):

factor \[x^2-2x -3=0 \]

OpenStudy (raden):

note for this case : x must be a positive number (x > 0)

OpenStudy (anonymous):

Actually \[ x\ge -\frac 32 \]

OpenStudy (raden):

i see, but is it possible the left side would give the negative value (in right side) ??